OpenStudy (anonymous):

estimate 3 2/5 / 2 2/7

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):



OpenStudy (anonymous):

17/5 / 16/7 17/5 * 7/16 = 119/80

OpenStudy (anonymous):

estimate & find the act

OpenStudy (anonymous):

moyo im as confused as you are lol

OpenStudy (anonymous):

haha. how do we know it's wrong neenee?

OpenStudy (anonymous):

30/17

OpenStudy (anonymous):

for the secound one

OpenStudy (anonymous):

for the first one i still got 119/80

OpenStudy (anonymous):

cuz when i put the answers on my mymathlab where i do all my hw it says its wrong

OpenStudy (anonymous):

15/2 - 17/4 = 30/4 - 17/4 = 13/4

OpenStudy (anonymous):

the mathlab s wrong then

OpenStudy (anonymous):

did you write out the problem correctly? we both got 119/80. what answer does mathlab give you?

OpenStudy (anonymous):

it gave me 2 2/9 something like that

OpenStudy (anonymous):

its okk ill just leave it

OpenStudy (anonymous):

oh..i get what it wants

OpenStudy (anonymous):

hold on i think i can figure it out

OpenStudy (anonymous):

no wait it was 3/2

OpenStudy (anonymous):

the answer was 3/2?

OpenStudy (anonymous):

heres another 1

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

f this im out

OpenStudy (anonymous):

jk ask another

OpenStudy (anonymous):

lol its ok forget it calm it man

OpenStudy (anonymous):

\[1\frac{39}{80} \]

OpenStudy (anonymous):

1/9-5/6 (10)